OpenStudy (anonymous):

Jaxon purchased his first home for $155,600. He obtained a 30-year, fixed-rate mortgage with an interest rate of 7.15% in the amount of $150,400. The seller paid $2,180.45 in property taxes for the coming year. If the date of the closing was June 22, and Jaxon owned the property on that date, how much will he owe in prorated taxes and interest? • $265.16 • $1,292.00 • $1,152.21 • $1,417.35

OpenStudy (phi):

I would think he has to pay for the taxes from June 22 to the end of the year June 22 is 173 day of the year, so he owes 365-173+1 days out of 365 or 193/365 of the taxes I think he has to pay the interest on the loan up to the end of the month june 22 through june 30, or 9 days.

OpenStudy (anonymous):

ok. so I would do 2180.45*(193/365) but does it mean total interest or just interest for that month?

OpenStudy (phi):

I think it is just the interest for 9 days

OpenStudy (anonymous):

How do I find the interest for just nine days?

OpenStudy (phi):

I think you divide the rate 0.0715 by 360 to get a daily rate.

OpenStudy (anonymous):

ok. so it would be 150400(((1+(.0715/360))^(9)-1)/(.0715/360)?

OpenStudy (anonymous):

No.. That is not right because then you get 1.4*10^6

OpenStudy (anonymous):

I guess that could be right and the answer is 1152.21

OpenStudy (phi):

9 days * 0.0175/360 * 150,400 (I am guessing) is close.

OpenStudy (anonymous):

so 1152.95+65.8= 1218.75

OpenStudy (anonymous):

How do you figure 65.8 is only for 1 day.. it seems like .0175/360*150400=7.31 would be for 1 day then we did 7.31*9 and got 65.8

OpenStudy (phi):

I was confused. You are correct unfortunately, none of the answers matches.

OpenStudy (anonymous):

yeah.. ill ask my teacher. thank you!

OpenStudy (phi):

oops, I see you used 0.0175 and it should be 0.0715

OpenStudy (phi):

I get interest for 9 days is 9*0.0715/360 * 150400= 268.84 that plus the taxes 1152.95 gives 1421.79 which is close to 1417.35 I would still ask your teacher, because it would be nice to get an exact match.